OpenStudy (anonymous):

P varies directly as q and inversely as y. One set of values is P = 0.064, q = 16 and y = 5. Find P when q = 12 and y = 10. P= Does anyone know how to do this?

OpenStudy (anonymous):

P = kq q = m/y

OpenStudy (anonymous):

Do you know what the number answer would be?

OpenStudy (anonymous):

You don't want to solve it?

OpenStudy (anonymous):

I don't understand how, or I would.

OpenStudy (anonymous):

See, P = kq You have P = 0.064, q = 16 Can you find k by plugging in the values?

OpenStudy (anonymous):

You would multiply 0.064, and 16?

OpenStudy (anonymous):

or divide?

OpenStudy (anonymous):

you should divide them.

OpenStudy (anonymous):

So it's .004?

OpenStudy (anonymous):

Yes you are right.. so .004 is the value of what??

OpenStudy (anonymous):

k?

OpenStudy (anonymous):

Yes.. Now second equation is: q = m/y q = 16 and y = 5 solve for m now..

OpenStudy (anonymous):

So you would multiply 16 and 5, to get 80.?

OpenStudy (anonymous):

See we have dong wrong.. Let us try once more..

OpenStudy (anonymous):

Okay, where did we go wrong?

OpenStudy (anonymous):

See, P varies directly with q so, P = kq Now you know P = 0.064, q = 16 Now find k.. Till here we were right.. What is the value of k you got?

OpenStudy (anonymous):

0.04?

OpenStudy (anonymous):

In earlier case you have got different.. check your answer once again..

OpenStudy (anonymous):

I'm sorry, I aren't getting this. :/